1. The length of a buffet table is 4 more than 3 times the width. If the perimeter is 40 feet, find the dimensions of the table.
2. One day Tom, Bob, and Jane sat down to figure out who was the oldest. Their combined ages came to 54. Tom is four times as old as Bob less 10. Bob is four less than Jane. Who is the oldest, and how old are they?

Answer by Mike(39)

1. Let the length = l and the width = w. Therefore:
l = 3w + 4
I assume the table is rectangular.
2l + 2w = 40
2(3w + 4) + 2w = 40
6w + 8 + 2w = 40
8w = 32
w = 4
l = 16
2. Sent you the answer for that via e-mail.